OpenStudy (anonymous):

Five less than twice a number is four more then three times the number. How do I write this down with numbers and stuff?

OpenStudy (anonymous):

Don't give up!

OpenStudy (anonymous):

x2 - 5 = 4 - 3x ? i think this is right.

OpenStudy (anonymous):

idk, I got 2 - 5 = 3 + 4, but i think yours make more sense